OpenStudy (anonymous):

use the substitution method to solve the linear system y=x+2 2x+y=8

OpenStudy (anonymous):

so for this one u can can substitute x+2 for y in the second equation: 2x+(x+2)=8

OpenStudy (anonymous):

now u solve it 3x+2=8 3x=6 x=2

OpenStudy (anonymous):

now u can use the value of x to find y y=2+2 y=4

OpenStudy (anonymous):

did u get it @HisWifey<3

OpenStudy (anonymous):

yea thnx yuh !